OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 2) =< 2?

OpenStudy (zzr0ck3r):

"which of the following" ?

OpenStudy (anonymous):

A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 x x 3

OpenStudy (texaschic101):

2(x - 2) <= 2 2x - 4 <= 2 2x <= 2 + 4 2x < = 6 x <= 3 a number line with a closed circle and shaded to the left (correct) The inequality signs did not show up.....x <= 3 or 3 >=x are both correct.